Mathematics 244 – Linear Algebra Definitions Quiz 6 – March 16, 2007

Your Name:

A) (5) Define: T : R^{n} → R^{m} is a linear transformation.

B) (5) Define: {v_{1}, …, v_{k}} is a linearly dependent set of vectors in R^{n}.